# Dissecting CLIP: Decomposition with a Schur Complement-based Approach

### v1

@@ A.2 Conditional Entropy Interpretation of SCE Score
Note that the above provides our definition of the Schur-Complement-Entropy for the image part $\mathrm{SCE_{I}}$:
@@ Appendix B Additional Individual Image Decomposition Results via SC-Based Method
Results for synthetic images generated using SDXL are shown in Figures 9 and 10. Corresponding results for DALL-E 3 are presented in Figures 11 and 12. Notably, the Schur Complement-based decomposition successfully isolates and removes image directions corresponding to a text condition that describes the concept to be excluded.
@@ Appendix C Additional Results on Diversity Evaluation
To further validate the findings presented in the main text, we conducted a similar experiment (Figure 6) using the Cosine Similarity Kernel. The results confirm that the diversity trends observed in the main text persist under a finite-dimensional kernel. Figures 13 and 14 illustrate the variation in SCE diversity when conditioned on different text prompts.
@@
Appendix D Additional Experiments on the Image Captioning Task
In the main text, we discussed SC-based decomposition for text-to-image models and demonstrated how images can be decomposed given text prompts. Here, we show that the reverse process is also possible. Specifically, we explored decomposing captions based on their corresponding images.
The experimental setup mirrors that of Figure 4, with the key difference being that, instead of generating images for text prompts, we generated captions for the corresponding images. For this task, we used gpt4o-mini as the captioning model. Figure 15 illustrates the experimental setup.
We selected images closely aligned with the concept we aimed to remove. For instance, to eliminate the "cat" direction in text, we used an image of a cat against a white background to better isolate the concept. After applying corrections for "animals" or "objects" in the text prompt, we observed successful decomposition, as reflected in the second column: the corrected CLIP embedding is no longer sensitive to the removed concept.
These findings highlight the versatility of the SCE method, demonstrating its applicability across a wide range of tasks that rely on a shared embedding space.
@@ Appendix D Additional Experiments on the Image Captioning Task
Appendix E Results with Naive Text Embedding Subtraction without considering the adjust matrix $\Gamma^{*}$
@@
Figure 17: Effect of removing encoded "cassette player" text on top of ImageNet sampled. Top figure represents naive method and bottom figure represents SC method.
@@
Appendix F Additional Numerical Results

### v2

@@ A.2 Conditional Entropy Interpretation of SCE Score
Note that the above provides our definition of the Schur-Complement-Entropy for the image part $\mathrm{SCE_{I}}$ and the text part $\mathrm{SCE_{T}}$ as follows:
@@
where $\lambda^{(\Lambda_{I})}_{j}$ denotes the $j$th eigenvalue of matrix $\Lambda_{I}$ and $\mathrm{Tr}(\Lambda_{I})=\sum_{j=1}^{d}\lambda^{(\Lambda_{I})}_{j}$ is the sum of the eigenvalues. Note that we follow the same definition for the text part $\Lambda_{T}$.
@@ Appendix B Additional Individual Image Decomposition Results via SC-Based Method
Results for synthetic images generated using SDXL are shown in Figures 12 and 13. Corresponding results for DALL-E 3 are presented in Figures 14 and 15. Notably, the Schur Complement-based decomposition successfully isolates and removes image directions corresponding to a text condition that describes the concept to be excluded.
To expand on the results in Figure 8, we constructed a dataset of animals with traffic signs using FLUX.1-schnell [27]. Figure 20 illustrates that after canceling either of the subjects using the Schur complement method, Kernel-PCA clusters according to the remaining concepts in the image.
@@ Appendix C Additional Results on Diversity Evaluation
To further validate the findings presented in the main text, we conducted a similar experiment (Figure 5) using the Cosine Similarity Kernel. The results confirm that the diversity trends observed in the main text persist under a finite-dimensional kernel. Figures 16 and 17 illustrate the variation in SCE diversity when conditioned on different text prompts.
Similar to Figure 5, we conducted similar experiment with animals and objects dataset in Figure 18. Our resuts mirror previous findings, strengthening the proposed diversity evaluation metric in measuring subject quantity related diversity.
Moreover, we evaluate the diversity of typographically attacked ImageNet samples in Figure 21. Specifically, we overlay the text "cassette player" onto images from 10 different ImageNet classes and measure diversity as the number of distinct classes increases. The presence of overlaid text diverts CLIP’s sensitivity away from image content, causing it to encode the direction indicated by the text instead.
To illustrate this effect, we visualize salience maps of CLIP embeddings given a prompt referring to an object behind the overlayed text. The results show that CLIP is highly sensitive to centrally placed text, even when it is unrelated to the prompted object. However, applying SC-based decomposition mitigates this bias. This correction is reflected in the diversity plots, where $SCE_{I}$ increases rapidly as the number of ImageNet classes grows, whereas Vendi, Coverage, and Recall exhibit much weaker correlations with class diversity.
@@
Figure 18: Evaluated SCE and Vendi scores with Gaussian Kernel on different animals with objects.
@@
Figure 21: Evaluated SCE, Vendi, Recall and Coverage scores with Gaussian Kernel on ImageNet with overlayed text.
Appendix D Results with Naive Text Embedding Subtraction without considering the adjust matrix $\Gamma^{*}$
@@
Figure 23: Effect of removing encoded "cassette player" text on top of ImageNet samples. Top figure represents naive method and bottom figure represents SC method.
@@
Appendix E Additional Numerical Results